OpenStudy (anonymous):

how do i solve this? 6x(x-1)2 (to the second power)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[6x(x-1)^{2}\]

OpenStudy (anonymous):

X= -2

OpenStudy (anonymous):

this is not an equation. by "solve ' i take it you mean "expand"

OpenStudy (anonymous):

\[6x(x-1)^2=6x(x-1)(x-1)=6x(x^2-2x+1)=6x^3-12x^2+6x\]